OpenStudy (anonymous):

The manager of a candy shop sells chocolate covered peanuts for $8 per pound and chocolate covered cashews for $15 per pound. The manager wishes to mix 80 pounds of the cashews to get a mixture that will sell for $9 per pound. How many pounds of peanuts should be used?

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

OK either I am not figuring this out right or just lost it since it is Monday :( The choices we are given are 280, 240, 560, and 480 but when I figure the amount of peanuts I get 549 rounded off???

OpenStudy (anonymous):

*number you want

OpenStudy (anonymous):

thats what I did. which gave me 11.429 if divided and then subtract from 80 and multiply by 80?

OpenStudy (anonymous):

so about 11.4 pounds of cashew and the rest, 68.6 pounds of peanuts

OpenStudy (anonymous):

multiply by 8 sorry

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Sorry word problems really confuse me

OpenStudy (anonymous):

i screwed this all up because i read it wrong. i thought the total was 80 pounds but it is 80 pounds of cashews. so we have to start from the beginning every thing i wrote is wrong

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

we have 80 pounds of cashews, they are worth \(80\times 15=\$1200\)

OpenStudy (anonymous):

you add \(x\) pounds of peanuts, they are worth \(8\times x\) so the total cost is \[1200+8x\] and the total number of pounds is \(80+x\) which should cost $9 a pound, so we solve \[1200+8x=9(80+x)\] for x

OpenStudy (anonymous):

480

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

this should work \[1200+8x=720+9x\] \[1200-720=9x-8x\] \[x=480\]

OpenStudy (anonymous):

yeah it was better the second time. i jumped the gun sorry

OpenStudy (anonymous):

Its ok huge help tho!

OpenStudy (anonymous):

Thanks agaiN!

OpenStudy (anonymous):

yw for real this time